1.     Suppose that f is defined recursively by:

     f (0) 5= and f n( + =1) 2fn+5. Find f(1), f(2), f(3) and f(4)? 


1
Expert's answer
2021-04-27T13:32:34-0400

We have given that,

"f(0) = 5"

"f(n+1) = 2f(n)+5"


Putting, "n=0"

"f(1) = 2f(0)+5 = 2 \times 5+5 = 15"


Putting, "n = 1"

"f(2) = 2f(1)+5 = 2 \times 15 +5 = 35"


Putting, "n=2"

"f(3) = 2f(2)+5 = 2 \times 35+5 = 75"


Putting "n = 3"

"f(4) = 2f(3)+5 = 2 \times 75 +5 = 155"